use squeeze theorem to show that: limit as x approaches infinity of (sin(e^x))/x =0, hense evaluate the limit as x approaches infinity of (sin(e^x))/sqr(x^2+2)
1
Expert's answer
2021-07-19T05:50:08-0400

Suppose "x\not=0."


"0\leq|\dfrac{\sin(e^x)}{x}|\leq \dfrac{1}{|x|}, x\in\R, x\not=0"

"\lim\limits_{x\to\infin}\dfrac{1}{|x|}=0"

Then by the Squeeze Theorem


"\lim\limits_{x\to\infin}|\dfrac{\sin(e^x)}{x}|=0"

"0\leq|\dfrac{\sin(e^x)}{\sqrt{x^2+2}}|\leq |\dfrac{\sin(e^x)}{x}|, x\in\R, x\not=0"

Then by the Squeeze Theorem


"\lim\limits_{x\to\infin}|\dfrac{\sin(e^x)}{\sqrt{x^2+2}}|=0"

Therefore


"\lim\limits_{x\to\infin}\dfrac{\sin(e^x)}{\sqrt{x^2+2}}=0"